OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4

OpenStudy (zzr0ck3r):

distribute first 3x+x = -2x+2-4 then collect like terms 3x+x+2x = 2-4 then combine 6x = -2 solve for x x = -2/6 = -1/3